Please show me how to solve this problem
find f(-3) if f(x)= 2x-1

Answer by rapaljer(4667)

f(x) = 2x - 1

First, this is a special notation that does NOT mean f times x. It is a special notation that is read "f of x", and it means that you are to substitute a value of x into this formula and simplify the result to find the value of the answer. It is MUCH easier than everyone can possibly imagine.
Given
f(x) = 2x - 1
f(-3) means that you sustitute the value of -3 in place of the x value(s).
f(-3) = 2(-3) - 1
f(-3) = -6 -1 = -7

You didn't ask for it, but if you need another example, suppose you wanted to find f(9) or f(-7).
f(9) = 2(9) - 1 = 18-1 = 17

f(-7) = 2(-7) - 1 = -14 -1 = -15

R^2 at SCC